Fermi National Laboratory


Accelerator Update for December 3, 2004 - December 6, 2004

Friday December 3
The day shift began with the Antiproton Source (Pbar) stacking, with the Tevatron (TeV), Recycler, Main Injector (MI), and Switchyard (SY) conducting studies, and with MiniBooNE taking beam.

Operations receive permission to send beam to NuMI at 8:21 AM.

Operators had to turn off the TeV at 10 AM to let technicians to work on some quadrupoles (BQ2-4). Experimenters at D-Zero accessed their collision hall at the same time.

At 1:45 PM, NuMI experts began taking beam and working it down their beamline. By 6:50 PM, NuMI experts had beam to their absorber. At 9:04 PM, on their tenth pulse of beam, NuMI experts declared their tuning a success; they had gotten beam to their hadron detector. Below is the multiwire profile of their beam.

NuMI Line Profiles v5.4

Operations turned off the NuMI beam at 9:54 PM.

Operations reported at midnight that SY experts were having trouble getting beam past TeV sector F44.

Saturday December 4
The midnight shift began with operations monitoring the Pbar stack, and with Recycler, SY, and MiniBooNE conducting studies.

Pbar resumed stacking at 12:55 AM.

Operations started injecting protons into the TeV for a damper study at 1:30 AM.

At 4:44 AM, operators reported having problems with Pbar emittances.

NuMI experts began a study period at 8:56 AM.

At 10:08 AM, the TeV ramp dumped due to a sector D4 lead failure.

At 10:09 AM, a Pbar Accumulator quadrupole power supply failed. Pbar lost its stack. Operators called in power supply experts to investigate.

At 11:10 AM, the TeV ramp dumped again due to a D4 lead failure.

At noon, operators reported seeing a drop in the output of the H- Source. Operators investigated.

At 1:55 PM, the TeV D4 lead failed again.

A Preaccelerator expert switched to the I- Source.

Operators prepared the TeV for an access at 3:58 PM so that experts could investigate the sector D4 spool piece.

Pbar resumed stacking at 5:05 PM. Experts had found a bad ribbon cable for A:LQ.

Operations turned off MI and prepared TeV sector F for an access at 5:32 PM. SY experts searched for a beam pipe obstruction.

SY experts came out of access and MI was back on by 7:08 PM. MiniBooNE had beam and Pbar resumed stacking by 7:24 PM.

TeV began a 980 GeV ramp at 8:28 PM, but then tripped off due to a D3 lead failure.

NuMI experts finished their studies at 9:08 PM.

SY experts reported at 9:13 PM that they couldn't get beam beyond TeV sector F39.

Pbar expert reported at 11:44 PM that they found some bad equipment associated with their damper system, which was causing their emittance problems.

Sunday December 5
The midnight shift began with Pbar and TeV conducting studies, and with MiniBooNE taking beam.

A Pbar expert finished his study at 3:48 AM.

The TeV suffered a quench at sector F3. Beam was accidentally injected on top of beam.

Operations had established a QBS ramp at 7:05 AM, but the TeV quenched at F1, 2, & 3 while performing the QBS test.

Operations established a 980 GeV ramp in the TeV at 7:45 AM.

Around 8 AM Pbar resumed reverse proton studies.

At 10:10 AM, the TeV beam aborted due to lead failure at D0.

At 3:40 PM, the TeV suffered a multi-house quench.

Pbar began stacking at 4:48 PM

At 6:16 PM, the TeV ramp tripped off due to E2 lead failure. Cryo system techs replaced the E2 orifice at 9:01 PM.

A Recycler expert began a study period at 9:01 PM.

At 9:28 PM the TeV ramp dumped due to an E2 lead failure. Cryo system techs replaced the E2 & 3 orifices at 10:33 PM.

Monday December 6
The midnight shift began with Pbar stacking, with the TeV on a QBS ramp, with Recycler conducting studies, and with MiniBooNE taking beam.

At 12:14 AM, the TeV quenched at C1 and D4. Experts investigated. The TeV resumed ramping at 3:23 AM. Operations injected protons into the TeV at 4:30 AM and a Tevatron Electron Lens (TEL) expert began a study period. He declared the TEL operational at 5:08 AM.

Plans for Monday
The plan for today is to shoot to the TeV and create a small store for study purposes.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/6/2004   email Fermilab